OpenStudy (eric_d):

Express Partial fraction 2x^2+3x+2/x^2+3x+2 @Luigi0210

OpenStudy (luigi0210):

Alright, do you remember what we did on the last one?

OpenStudy (eric_d):

yes

OpenStudy (luigi0210):

Notice how they have the same powers, start by simplifying that first

OpenStudy (eric_d):

How to get 2 from this questions

OpenStudy (luigi0210):

2 as the leading constant?

OpenStudy (eric_d):

Nvm, I'll simplify it first

OpenStudy (eric_d):

x^2+3x+2= (x+1)(x+2) It can also be =(x+2)(x+1) Will this difference affect the final answer

OpenStudy (anonymous):

no..there no restriction for arrangement of terms..you can alternate them...

OpenStudy (luigi0210):

\[\Large \frac{2x^2+3x+2}{x^2+3x+2}=\frac{2x^2}{x^+3x+2}+\frac{3x+2}{x^2+3x+2}\] \[\Large =2+(\frac{3x+2}{(x+2)(x+1)})\]

OpenStudy (luigi0210):

That should be x^2+3x+2, not x+3x+2, Whoops

hartnn (hartnn):

thats not correct, please do long division

hartnn (hartnn):

the method itself is not correct

OpenStudy (eric_d):

So, long division is the only method to get 2

OpenStudy (eric_d):

@hartnn

hartnn (hartnn):

long divsion or adjusting the numerator (which does not work always) here, long division is the only option

OpenStudy (eric_d):

okay

hartnn (hartnn):

if co-efficients of highest power of x in num. and denom. are equal, only then , adjusting the numerator will work .

hartnn (hartnn):

example : Num = x^2 +ax+b Denom = x^2 +cx+d Then we express Num. as [x^2 + cx+d] - (ax+b) and then separate the numerator

OpenStudy (eric_d):

ok

OpenStudy (eric_d):

x^2/(x+1)(x+2)= A/(x+1)+ B(x+2) x^2/(x+2)(x+1)=A(x+2) + B(x+1) Are these considered the same Will it affect the final answer?

hartnn (hartnn):

why just x^2 ?

OpenStudy (eric_d):

Sorry, I don't get you

hartnn (hartnn):

lets try adjusting the numerator here \(2x^2+3x+2 = 2 (x^2+3x+2) - (3x-2)\) now you can separate out the numerator

hartnn (hartnn):

\(2x^2+3x+2 = 2 (x^2+3x+2) - (3x+2)**\)

hartnn (hartnn):

2- (3x+2)/(x^2+3x+2) now you can apply partial fractions method on (3x+2)/(x^2+3x+2)

OpenStudy (eric_d):

 @hartnn

hartnn (hartnn):

that is denominator i expressed numerator in terms of denominator

OpenStudy (eric_d):

Okay

OpenStudy (eric_d):



hartnn (hartnn):

let me give you a general example : if numerator = Ax^2 +Bx+C and denominator = Dx^2 +Ex+F then i express num = Ax^2 +Bx+C as A/D(Dx^2 +Ex+F) - (AE/Dx +AF/D) + Bx+C it may seem not so easy, but with specific examples and some practice, you will get hold of it

hartnn (hartnn):

let me take some numbers N = 3x^2 +4x+5 D = x^2 +9x+10 so i will express N as 3 (x^2 +9x+10) - (23x+ 25)